How many moles are in 4.1x10 to the 22nd power molecules of NaCl


1
Expert's answer
2021-03-03T01:54:38-0500

Q168125

How many moles are in 4.1x10 to the 22nd power molecules of NaCl?


Solution:


We are given the number of molecules of NaCl and have to convert it to 'moles of NaCl'.


One mole of a substance contains Avagadro's number of molecules or atoms.


1 mol = 6.022 * 1023 molecules.


Using this as a conversion factor, we can find the '# of moles of NaCl' in 4.1 * 1022 molecules of NaCl.



"No.\space of\space moles\space of\space NaCl 
= 4.1 * 10^{22} \space molecules\space of\space NaCl * \frac{1\space mol\space NaCl }{6.022 * 10^{23}\space molecules\space of\space NaCl }"


"= 0.0681\space mol\space of\space NaCl"




In the question, we are given 4.1 * 1022 molecules in 2 significant figures, so our final answer must also be in 2 significant figures.


Hence there are 0.068 moles of NaCl in 4.1 * 1022 molecules of NaCl.


Sometimes you may have to submit your answer in scientific notation.


Move the decimal point two times to the right to get 6.8 * 10-2 moles of NaCl as the answer.


The power of 10 is decreased by 2 because we have moved the decimal point two times to the right side.


Also, refer to the image I have uploaded.





Thank you. :)